OpenStudy (anonymous):

1. If you have 1.50 Liters of solution that contains 62.5 grams of magnesium chloride. What is the Molarity of that solution? Please show all work for any credit. 2. You have .62 L of a 2.0 M HCl Solution. How many moles of HCl do you have? How many grams of HCl? Please show all work for any credit.

OpenStudy (abb0t):

What happens if I don't show work, do I not get credit for your assignment? :(

OpenStudy (aaronq):

^LOL

OpenStudy (anonymous):

You seem to have no idea of how to use the simplest formula Do you know the formula to find concentration? C = n / V where n = moles and V = volume in L??? Well manipulat eht formula to find n. n = C x V. Why you would work on this all day is beyond lme. You did absolutely no research obviously. For the second problem you need to know another EXTREMELY SIMPLE formula. n = m / M moles = mass / molar mass so convert the 62.5 g of MgCL2 into moles then use the first formula above to find C.

OpenStudy (anonymous):

google helps most questions can be found there

OpenStudy (anonymous):

1) Knowing this, the first problem can be solved by multiplying .62 L by the 2.0 M HCl, this is because 2.0 M=x/.62 L, where x is the number of moles of solute. Therefore, just solve for x, algebraically. 2) First, convert the gram value of magnesium chloride to moles by dividing 62.5 g by the molar mass of magnesium chloride, this is found by just adding the molar masses of magnesium to chloride.. Therefore, 62.5 g MgCl x (1 mole/Molar Mass of MgCl). Once you have the molar value of MgCl, just divide it by the given 1.50 L to find molarity.